Which amino acid has a net charge of “+ 0.09” at pH 7? Show all calculations. Draw the structure of the amino acid that would predominate at this pH. (The answer should be Histindine; draw the structure of histidine. Calculate the net charge using a H-H equation. Please walk through every single step as if I was five years old I am very confused).
Answer #1

Henderson Hasselbach equation

pH = pKa + log (base/acid)

Base / acid = 10 ^ (pH - pKa)

Base + acid =1 (i.e. total number of acid and base molecules is 1)

So base = 1- acid

So , 1-acid / acid = 10^ (pH- pKa)

Base = 1/ 1+(10^ (pH-pKa))

Acid = 1/1+(10^(pKa-pH))

Base + Acid =0.09

pH=7

Putting the numbers

PKa =6.1 ,i.e. histidine

At this pH protonated form predominates.